# Convergence rates for polynomial optimization on set products

### v1

@@ Abstract
We consider polynomial optimization problems on Cartesian products of basic compact semialgebraic sets. The solution of such problems can be approximated as closely as desired by hierarchies of semidefinite programming relaxations, based on classical sums of squares certificates due to Putinar and Schmüdgen. When the feasible set is the bi-sphere, i.e., the Cartesian product of two unit spheres, we show that the hierarchies based on the Schmüdgen-type certificates converge to the global minimum of the objective polynomial at a rate in $O(1/t^{2})$, where $t$ is the relaxation order. Our proof is based on the polynomial kernel method. We extend this result to arbitary sphere products and give a general receipe to obtain convergence rates for polynomial optimization over products of distinct sets. Eventually, we rely on our results for the bi-sphere to analyze the speed of convergence of a semidefinite programming hierarchy approximating the order $2$ quantum Wasserstein distance.
@@ 4.1 Problem statement and main result
Given $n\in\mathbb{N}$, let $\rho,\nu\in\mathbb{C}^{n\times n}$ be two normalized quantum states, i.e., Hermitian matrices with nonnegative eigenvalues and unit trace. A quantum transport plan between $\rho$ and $\nu$ is a finite set of triples $\{(\omega_{\ell},u_{\ell},v_{\ell})_{\ell}\}$ such that
@@
where $\omega_{\ell}>0$, $\sum_{\ell}\omega_{\ell}=1$, $u_{\ell},v_{\ell}\in\mathbb{C}^{n}$, $\|u_{\ell}\|=\|v_{\ell}\|=1$. The set of all such quantum transport plans between two states $\rho$ and $\nu$ is denoted by $\mathcal{Q}(\rho,\nu)$. Given a plan $e=\{(\omega_{\ell},u_{\ell},v_{\ell})_{\ell}\}$, the associated (order $2$) quantum transport cost is defined as
@@
In [GLS22, Section 5.2], the authors propose a moment reformulation of the DPS hierarchy [DPS02], initially designed to distinguish separable and entangled states. Inspired by this reformulation, it is proved in [CM25] that $W_{2}^{2}(\rho,\nu)$ is the optimal value of a particular instance of the generalized moment problem (GMP), i.e., an infinite-dimensional linear problem over positive Borel measures:
@@
It is proved in [CM25] that the corresponding sequence of lower bounds given by this hierarchy converges to the order $2$ quantum Wasserstein distance, i.e., $W^{2}_{2}(\rho,\sigma)_{t}\to W^{2}_{2}(\rho,\sigma)$ when $t$ goes to infinity. In the recent contribution [GM25], the authors analyze convergence rates of hierarchies of lower bounds for the GMP. We recall this result in the case where the GMP instance involves a single measure and equality constraints only.
@@
Our final result, given in Theorem 11 below, provides the convergence rate of the moment hierarchy (39) for the order $2$ quantum Wasserstein distance. This result is obtained by combining Lemma 10 with our analysis derived in Theorem 3.
@@
where $\kappa(n)$ is a constant depending only on $n$ and being proportional to the constant $C_{\mathbf{X}}(2n,4)$ from Theorem 3. See relation (49) for details.
The proof of Theorem 11 is postponed in Section 4.3.
@@ 4.2 Changing variables from complex to real
s.t. $f_{\operatorname{Re}}\geq\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\Gamma_{1}+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})\Lambda_{1}+(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})\Gamma_{2}$
$\quad\quad\quad+(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})\Lambda_{2})\text{ on }\mathbf{X}_{\operatorname{Re}},$
@@ 4.3 Proof of Theorem 11
    there exists $(\Lambda_{i},\Gamma_{i})$ such that $\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\Gamma_{1}+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})\Lambda_{1}+(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})\Gamma_{2}+(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})\Lambda_{2})$ is positive on $\mathbf{X}_{\operatorname{Re}}$;
@@
$\text{with }\kappa(n):=\frac{3}{2}C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}\left(3f_{\operatorname{Re},\max}+\|w^{\operatorname{opt}}\|_{1}h_{\max}\right),$
@@
$f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=\operatorname{Tr}[(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})\overline{(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})}]$ $=|\sum_{i=1}^{n}x_{i}^{2}|^{2}+|\sum_{i=1}^{n}y_{i}^{2}|^{2}-2|\sum_{i=1}^{n}x_{i}y_{i}|^{2}$
$\leq\sum_{i=1}^{n}|x_{i}|^{2}+\sum_{i=1}^{n}|y_{i}|^{2}.$
@@
2. Let us show that the second quantity $\|w^{\operatorname{opt}}\|_{1}$ is less than $4n$. Note that every tuple $(\Lambda_{1},\Lambda_{2},\Gamma_{1},\Gamma_{2})$ achieving the supremum in the real dual (48) corresponds to a tuple $(\Lambda,\Gamma)$ achieving the supremum for the complex dual (38), with $\Lambda=\Lambda_{1}+\mathbf{i}\,\Lambda_{2}$ and $\Gamma=\Gamma_{1}+\mathbf{i}\,\Gamma_{2}$.
One has $0\leq\operatorname{Tr}(\Lambda\Lambda^{*})=\operatorname{Tr}(\Lambda^{2})=\operatorname{Tr}(\Lambda_{1}^{2})-\operatorname{Tr}(\Lambda_{2}^{2})$. This fact combined with Cauchy-Schwartz inequality implies that
$\|\Lambda_{1}\|_{1}+\|\Lambda_{2}\|_{1}\leq\sqrt{n}\left(\sqrt{\operatorname{Tr}(\Lambda_{1}^{2})}+\sqrt{\operatorname{Tr}(\Lambda_{2}^{2})}\right)\leq 2\sqrt{n}\sqrt{\operatorname{Tr}(\Lambda^{2})}\leq 2n\lambda_{\max}(\Lambda),$
and similarly one gets $\|\Gamma_{1}\|_{1}+\|\Gamma_{2}\|_{1}\leq 2n\lambda_{\max}(\Gamma)$.
In addition the polynomial $f-\mathbf{x}^{*}\Lambda\mathbf{x}-\mathbf{y}^{*}\Gamma\mathbf{y}$ is nonnegative on $\mathbf{X}$, implying the following
$\lambda_{\max}(\Lambda)+\lambda_{\max}(\Gamma)=\max_{\mathbf{x}^{*}\mathbf{x}=1}\mathbf{x}^{*}\Lambda\mathbf{x}+\max_{\mathbf{y}*\mathbf{y}=1}\mathbf{y}^{*}\Gamma\mathbf{y}\leq f_{\max}=2.$
@@
$\kappa(n)\leq\frac{3}{2}(3+4n)C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}.$ (49)

### v2

@@ Abstract
We consider polynomial optimization problems on Cartesian products of basic compact semialgebraic sets. The solution of such problems can be approximated as closely as desired by hierarchies of semidefinite programming relaxations, based on classical sums of squares certificates due to Putinar and Schmüdgen. When the feasible set is the bi-sphere, i.e., the Cartesian product of two unit spheres, we show that the hierarchies based on the Schmüdgen-type certificates converge to the global minimum of the objective polynomial at a rate in $O(1/t^{2})$, where $t$ is the relaxation order. Our proof is based on the polynomial kernel method. We extend this result to arbitrary sphere products and give a general recipe to obtain convergence rates for polynomial optimization over products of distinct sets. Eventually, we rely on our results for the bi-sphere to analyze the speed of convergence of a semidefinite programming hierarchy approximating the order $2$ quantum Wasserstein distance.
@@ 4.1 Problem statement and main result
Given $n\in\mathbb{N}$, let $\rho,\nu\in\mathbb{C}^{n\times n}$ be two normalized quantum states, i.e., Hermitian matrices with nonnegative eigenvalues and unit trace. A quantum transport plan between $\rho$ and $\nu$ is a finite set of triples $\{(\omega_{\ell},u_{\ell},v_{\ell})\}_{\ell}$ such that
@@
where $\omega_{\ell}>0$, $\sum_{\ell}\omega_{\ell}=1$, $u_{\ell},v_{\ell}\in\mathbb{C}^{n}$, $\|u_{\ell}\|=\|v_{\ell}\|=1$. The set of all such quantum transport plans between two states $\rho$ and $\nu$ is denoted by $\mathcal{Q}(\rho,\nu)$. Given a plan $e=\{(\omega_{\ell},u_{\ell},v_{\ell})\}_{\ell}$, the associated (order $2$) quantum transport cost is defined as
@@
In [GLS22, Section 5.2], the authors propose a moment reformulation of the DPS hierarchy [DPS02], initially designed to distinguish separable and entangled states. Inspired by this reformulation, it is proved in [CM25, Theorem 8] that $W_{2}^{2}(\rho,\nu)$ is the optimal value of a particular instance of the generalized moment problem (GMP), i.e., an infinite-dimensional linear problem over positive Borel measures:
@@
It is proved in [CM25, Theorem 10] that the corresponding sequence of lower bounds given by this hierarchy converges to the order $2$ quantum Wasserstein distance, i.e., $W^{2}_{2}(\rho,\sigma)_{t}\to W^{2}_{2}(\rho,\sigma)$ when $t$ goes to infinity. In the recent contribution [GM25], the authors analyze convergence rates of hierarchies of lower bounds for the GMP. We recall this result in the case where the GMP instance involves a single measure and equality constraints only.
@@
Our final result, given in Theorem 12 below, provides the convergence rate of the moment hierarchy (39) for the order $2$ quantum Wasserstein distance, assuming that there is a bounded dual optimal solution of (38). This result is obtained by combining Lemma 10 with our analysis derived in Theorem 3. Let us denote by $\rho_{\min}$ and $\rho_{\max}$ the minimal and maximal respective eigenvalues of $\rho$, respectively. We define similarly $\nu_{\min}$ and $\nu_{\max}$. To bound the dual solution of (38), we will assume the following.
Assumption 11.
Both $\rho$ and $\nu$ have full rank and either $\nu_{\min}>\rho_{\max}$ or $\rho_{\min}>\nu_{\max}$.
@@
where $\kappa(n)$ is a constant depending only on $n$, $\rho$ and $\nu$, and being proportional to the constant $C_{\mathbf{X}}(2n,4)$ from Theorem 3. See relation (50) for details.
The proof of Theorem 12 is postponed in Section 4.3. Note that the hypotheses of Assumption 11 on $\rho$ and $\nu$ are rather restrictive. One can instead simply assume that there is an optimal solution $w^{\operatorname{opt}}=(\Lambda,\Gamma)$ for the dual (38) with bounded entries, in which case the constant $\kappa(n)$ is given as in (4.3).
@@ 4.2 Changing variables from complex to real
s.t. $f_{\operatorname{Re}}\geq\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\Lambda_{1}+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})\Gamma_{1}+(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})\Lambda_{2}$
$\quad\quad\quad+(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})\Gamma_{2})\text{ on }\mathbf{X}_{\operatorname{Re}},$
@@ 4.3 Proof of Theorem 12
    there exists $(\Lambda_{i},\Gamma_{i})$ such that $\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\Lambda_{1}+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})\Gamma_{1}+(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})\Lambda_{2}+(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})\Gamma_{2})$ is positive on $\mathbf{X}_{\operatorname{Re}}$;
@@
$\text{with }\kappa(n):=\frac{3}{2}C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}\left(3f_{\operatorname{Re},\max}+\|w^{\operatorname{opt}}\|_{1}h_{\max}\right),$ (49)
@@
$f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=\operatorname{Tr}[(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})\overline{(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})}]$ $=\left|\sum_{i=1}^{n}x_{i}^{2}\right|^{2}+\left|\sum_{i=1}^{n}y_{i}^{2}\right|^{2}-2\left|\sum_{i=1}^{n}x_{i}y_{i}\right|^{2}$
$\leq\left(\sum_{i=1}^{n}|x_{i}|^{2}\right)^{2}+\left(\sum_{i=1}^{n}|y_{i}|^{2}\right)^{2}.$
@@
2. Under Assumption 11, let us handle specifically the case where $\nu_{\min}>\rho_{\max}$, the other case can be treated analogously. Let us show that in this case the second quantity $\|w^{\operatorname{opt}}\|_{1}$ is less than $2\sqrt{2}n\left(1-\frac{\rho_{\max}^{2}}{\nu_{\min}^{2}}\right)$. Note that every optimal tuple $(\Lambda_{1},\Lambda_{2},\Gamma_{1},\Gamma_{2})$ for the real dual (48) corresponds to an optimal tuple $(\Lambda,\Gamma)$ for the complex dual (38), with $\Lambda=\Lambda_{1}+\mathbf{i}\,\Lambda_{2}$ and $\Gamma=\Gamma_{1}+\mathbf{i}\,\Gamma_{2}$.
For all $z=a+\mathbf{i}\,b$ with $a,b\in\mathbb{R}$, one has $|a|+|b|\leq\sqrt{2}\sqrt{a^{2}+b^{2}}=\sqrt{2}|z|$, thus $\|\Lambda_{1}\|_{1}+\|\Lambda_{2}\|_{1}\leq\sqrt{2}\|\Lambda\|_{1}$, and $\|\Gamma_{1}\|_{1}+\|\Gamma_{2}\|_{1}\leq\sqrt{2}\|\Gamma\|_{1}$. For all $\delta>0$, the tuple $(\Lambda+\delta I,\Gamma-\delta I)$ is also optimal for (38) since
$\operatorname{Tr}(\rho(\Lambda+\delta I)+\nu(\Gamma-\delta I))=\operatorname{Tr}(\rho\Lambda+\nu\Gamma)+\delta\operatorname{Tr}(\rho-\nu))=\operatorname{Tr}(\rho\Lambda+\nu\Gamma),$
and
$\mathbf{x}^{*}(\Lambda+\delta)I\mathbf{x}-\mathbf{y}^{*}(\Gamma-\delta I)\mathbf{y}=\mathbf{x}^{*}\Lambda\mathbf{x}-\mathbf{y}^{*}\Gamma\mathbf{y},$
for all $(\mathbf{x},\mathbf{y})\in\mathbf{X}$. Thus, we assume without loss of generality that $\Lambda,-\Gamma\succeq 0$. Let us note $\mathbf{e}_{i}$ the vector having only 0 components except a component 1 in the $i$-th place. Since the polynomial $f-\mathbf{x}^{*}\Lambda\mathbf{x}-\mathbf{y}^{*}\Gamma\mathbf{y}$ is nonnegative on $\mathbf{X}$, one has in particular for $\mathbf{x}=\mathbf{y}=\mathbf{e}_{i}$:
$\Lambda_{ii}+\Gamma_{ii}\leq f_{\max}=2,$
implying that $\operatorname{Tr}(\Lambda)\leq 2-\operatorname{Tr}(\Gamma)$. In addition, the objective function $\operatorname{Tr}(\rho\Lambda)+\operatorname{Tr}(\nu\Gamma)$ is always nonnegative, thus
$\rho_{\max}\operatorname{Tr}(\Lambda)\geq\operatorname{Tr}(\rho\Lambda)\geq-\operatorname{Tr}(\nu\Gamma)\geq-\nu_{\min}\operatorname{Tr}(\Gamma).$
After combining the two above inequalities we obtain $\operatorname{Tr}(\Lambda)-\operatorname{Tr}(\Gamma)\leq 2\left(1-\frac{\rho_{\max}^{2}}{\nu_{\min}^{2}}\right)$.
The desired bound follows then from the Cauchy-Schwartz inequality:
$\|\Lambda_{1}\|_{1}+\|\Lambda_{2}\|_{1}+\|\Gamma_{1}\|_{1}+\|\Gamma_{2}\|_{1}$ $\leq\sqrt{2}\left(\|\Lambda\|_{1}+\|\Gamma\|_{1}\right)$
$\leq\sqrt{2}\sqrt{n}\left(\sqrt{\operatorname{Tr}(\Lambda^{2})}+\sqrt{\operatorname{Tr}(\Gamma^{2})}\right)$
$\leq\sqrt{2}n(\Lambda_{\max}-\Gamma_{\min})\leq\sqrt{2}n(\operatorname{Tr}(\Lambda)-\operatorname{Tr}(\Gamma)).$
where $\Lambda_{\max}$ and $\Gamma_{\min}$ denotes the largest and smallest eigenvalues of $\Lambda$ and $\Gamma$, respectively.
@@
$\kappa(n)\leq\frac{3}{2}\left(3+4\sqrt{2}n\left(1-\frac{\rho_{\max}^{2}}{\nu_{\min}^{2}}\right)\right)C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}.$ (50)